English

In a Simultaneous Throw of a Pair of Dice, Find the Probability of Getting Neither 9 Nor 1 1 as the Sum of the Numbers on the Faces

Advertisements
Advertisements

Question

In a simultaneous throw of a pair of dice, find the probability of getting neither 9 nor 1 1 as the sum of the numbers on the faces

Advertisements

Solution

In a throw of pair of dice, total no of possible outcomes = 36 (6 × 6) which are

(1, 1) (1, 2) (1, 3) (1, 4) (1, 5) (1, 6)

(2, 1) (2, 2) (2, 3) (2, 4) (2, 5) (2, 6)

(3, 1) (3, 2) (3, 3) (3, 4) (3, 5) (3, 6)

(4, 1) (4, 2) (4, 3) (4, 4) (4, 5) (4, 6)

(5, 1) (5, 2) (5, 3) (5, 4) (5, 5) (5, 6)

(6, 1) (6, 2) (6, 3) (6, 4) (6, 5) (6, 6)

`barE` ⟶ event of getting neither 9 nor 11 as the sum of numbers on faces

E ⟶ getting either 9 or 11 as the sum of no’s on faces

No. of favorable outcomes = 6 {(3, 6) (4, 5) (5, 4) (6, 3) (5, 6) (6, 5)}

Total no. of possible outcomes = 36

We know that, Probability P(E) =`"No.of favorable outcomes"/"Total no.of possible outcomes"`

P(E) =6/36 =1/6
P`(barE)` = 1 − 𝑃 = 1 − 1/6 =5/6

shaalaa.com
  Is there an error in this question or solution?
Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×